OpenStudy (anonymous):

quick question..11/a+8/2a=?

OpenStudy (anonymous):

In the current form, no solutions exist. Maybe you wrote down wrong? http://www.wolframalpha.com/input/?i=solve%2811%2Fa+-+8%2F2a+%3D+0%2Ca%29

OpenStudy (anonymous):

Ok, equals 7. That's better.

OpenStudy (anonymous):

no that's a qustion mark lol

OpenStudy (anonymous):

11/a - 8/2a = 7 Multiply both sides by a. 11 - 8/2 = 7a Divide through by 7. 11/7 - 4/7 = a 7/7 = a 1 = a Hope this helps.

OpenStudy (anonymous):

ha thanks but that's a question mark lol

OpenStudy (anonymous):

I think seanwalsh misread your equation; he is correct if the second term on the LHS is 8/2a, but not if it is 8/(2a), which is what your drawing looks like to me. In case it is the latter:\[\frac{11}{a}+\frac{8}{2a}=7\]now reduce the 8/2 to a 4\[\frac{11}{a}+\frac{4}{a}=7\]combine like terms\[\frac{15}{a}=7\]or\[a=\frac{15}{7}\]